OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 2y = 6 x + 3y = −1 −2 2 5 −5

OpenStudy (anonymous):

method or answer?

OpenStudy (anonymous):

2x + 2y = 6 x + 3y = -1 y =-1/3 - 1/3x 2x + 2(-1/3 - 1/3x)=6 2x - 2/3 - 2/3x=6 4/3x = 20/3 4x = 20 x = 5